Mastery Problem: Corporations: Organization, Stock Transactions, and Dividends

Pranks, Inc.

Pranks, Inc. is a manufacturer of joke and novelty products for perpetrators of practical jokes. The corporation has paid several cash dividends throughout Year 6, the current year. It is also declaring a stock dividend to its stockholders as the calendar year-end approaches. You’ve been brought in as a consultant to assist with this process, and also to help determine whether some missing information can be determined before the distribution of the stock dividend is made. The company has two classes of stock: common stock and cumulative preferred stock.

Number of common shares authorized 900,000
Number of common shares issued 750,000
Par value of common shares $20
Par value of cumulative preferred shares $30
Paid-in capital in excess of par-common stock $7,000,000
Paid-in capital in excess of par-preferred stock $0
Total retained earnings before the stock dividend is declared $33,500,000
No treasury share have been reissued.
Preferred Dividends Common Dividends
Year Total Cash
Dividends
Total Per Share Total Per Share
Year 1 30,000   30,000 0.20       0 0.00      
Year 2 54,000   54,000 0.36       0 0.00      
Year 3 105,000   51,000 0.34       54,000 0.09      
Year 4 135,000   45,000 0.3       90,000 0.15      
Year 5 153,000   45,000 0.3       108,000 0.18      
Year 6 225,000   45,000 0.3       180,000 0.3      

Stock Dividend

The company declared a 2% common stock dividend on December 1, and would like you to compute the following pieces of missing information. The market value of the common shares is $25 on December 1, and is $32 on the actual distribution date of the stock, December 31.

Fill in the missing information in the following table, using the information given and your work on the other panels. All “before” items are before the stock dividend was declared. All “after” items are after the stock dividend was declared and closing entries were recorded at the end of the year.

Total paid-in capital before the stock dividend $
Total retained earnings before the stock dividend
Total stockholders’ equity before the stock dividend $
Total paid-in capital after the stock dividend $
Total retained earnings after the stock dividend
Total stockholders’ equity after the stock dividend $
Answer #1
Outstanding stock:
Common dividend is declared on common stock outstanding
Year 6,Total common dividend=$ 180000
Year 6,Common dividend per share=0.30
Common stock outstanding=180000/0.30=600000 shares
Preferred stock outstanding=Total preferred dividends in Year 6/Preferred dividends per share in year 6=45000/0.30=150000 shares
Preferred dividend as a % of par=Preferred dividends per share in year 6/par value of cumulative preferred shares=0.30/30=.01=1%
Cumualtive preferred dividends paid in year 2 and 3
Since the preferred dividends per share is more for year 2 and 3
Stock dividend declared
When stock dividend is less than 20-25%, it is considered as small stock dividend
Stock dividend is declared on common stock outstanding
Common stock issued=600000*2%=12000 shares
Journal entry:
Date Debit Credit
Dec 1. Retained earnings (600000*2%*25) 300000
Common stock dividend distributable (600000*2%*20) 240000
Paid-in capital in excess of par-common stock 600000*2%*(25-20) 60000
(Stock dividend declared)
Dec 31. Common stock dividend distributable 240000
Common stock 240000
(Stock dividend issued)
Total paid-in capital before stock dividend:
$
Common stock (750000*20) 15000000
Preferred stock (150000*30) 4500000
Paid-in capital in excess of par-common stock 7000000
Paid-in capital in excess of par-preferred stock 0
Total paid-in capital 26500000
Note: the market value of common shares on Date of declaration is relevant and on the date of actual distribution is not relevant
Total retained earnings before the stock dividend=$ 33500000
Total stockholder's equity before the stock dividend=Total paid-in capital+Total retained earnings-Treasury stock
Number of shares in treasury stock=Number of common shares issued-Number of common shares outstanding=750000-600000=150000
Let's assume that treasury stock acquired at par value
Treasury stock in $=150000*20=$ 3000000
Total stockholder's equity before the stock dividend=26500000+33500000-3000000=$ 57000000
Total paid-in capital after stock dividend:
$
Common stock (750000*20)+240000 15240000
Preferred stock (150000*30) 4500000
Paid-in capital in excess of par-common stock (7000000+60000) 7060000
Paid-in capital in excess of par-preferred stock 0
Total paid-in capital 26800000
Total retained earnings after the stock dividend=33500000-300000=$ 33200000
Total stockholder's equity after the stock dividend=Total paid-in capital+Total retained earnings-Treasury stock=26800000+33200000-3000000=$ 57000000